# Learning What Reinforcement Learning Can’t: Interleaved Online Fine-Tuning for hardest Questions

### v1

@@ 3.1 Experimental Setting
Evaluation. For our evaluation, we primarily consider five widely used mathematical reasoning benchmarks: AIME 2024, AIME 2025, AMC Li et al. (2024), OlympiadBench He et al. (2024), and MATH500 Hendrycks et al. (2021). Due to the relatively small test sets for AIME 2024, AIME 2025, and AMC, we report avg@32 for these datasets. For OlympiadBench and MATH500, we use pass@1 as the evaluation metric. Since our reinforcement learning training is mainly centered on mathematical reasoning, we also assess generalization performance on MMLU-Pro Wang et al. (2024). To prevent data contamination, we shuffle the multiple-choice options.
@@
Qwen-Math | 5.1 | 3146 | 1.4 | 3518 | 22.6 | 2611 | 37.2 | 1811 | 19.0 | 2429 | 30.5 | 938 | 19.3 | 2408 |
Qwen-Math-Instruct | 9.7 | 16453 | 8.2 | 15715 | 39.9 | 11734 | 77.2 | 7834 | 34.1 | 12424 | 33.0 | 20919 | 33.7 | 14179 |
SimpleRL-Zero | 27.0 | - | 6.8 | - | 54.9 | - | 76.0 | - | 34.7 | - | 34.5 | - | 38.98 | - |
@@
LUFFY | 27.2 | 3120 | 18.1 | 1852 | 61.1 | 1581 | 85.6 | 1370 | 53.6 | 1597 | 52.6 | 1052 | 49.7 | 1762 |
RL | 22.1 | 1131 | 13.9 | 1527 | 59.5 | 449 | 84.2 | 120 | 44.7 | 433 | 48.3 | 260 | 45.5 | 653 |
SFT | 26.9 | 16199 | 23.1 | 16609 | 58.9 | 9397 | 85.0 | 3717 | 50.7 | 9769 | 48.6 | 4906 | 48.9 | 10166 |
ReLIFT | 28.2 | 1666 | 20.1 | 1093 | 64.9 | 943 | 85.2 | 405 | 55.9 | 796 | 52.5 | 1463 | 51.1 | 1061 |
@@ 3.2 Main Results
Reasoning Performance. Our main results are presented in Table 1, where we compare ReLIFT with several representative zero-RL baselines, as well as direct SFT, on-policy RL, and off-policy RL method LUFFY Yan et al. (2025). Across five challenging competition-level reasoning benchmarks and one out-of-distribution benchmark, ReLIFT achieves the highest overall accuracy of 51.1%, outperforming all existing baselines and establishing a new state-of-the-art for models based on Qwen2.5-Math-7B. Specifically, ReLIFT exceeds the strongest zero RL baseline (Oat-Zero) by a substantial margin of +5.2 points in overall accuracy, and consistently achieves the best or second-best results on all individual benchmarks. Furthermore, compared to SFT and RL trained on the same dataset, ReLIFT demonstrates clear superiority: it outperforms SFT by +2.2 points and RL by +5.6 points in overall accuracy. ReLIFT also outperforms LUFFY, which combines on-policy and off-policy reasoning traces, despite using only 13% off-policy guidance. These results underscore the effectiveness of integrating RL with interleaved fine-tuning for hardest questions, enabling ReLIFT to achieve both higher performance and better generalization.
Reasoning Efficiency. In addition to accuracy, we evaluate the efficiency of different methods by measuring the average response length (LEN) across all benchmarks. ReLIFT demonstrates remarkable efficiency, generating shorter solutions (average LEN: 1,061) compared to SFT (average LEN: 10,166), Qwen2.5-Math-Instruct (average LEN: 14,179), and LUFFY, while still achieving superior accuracy. This indicates that ReLIFT not only produces more accurate answers but also does so with significantly less computational overhead and more concise reasoning chains.
@@ 3.3 Training Dynamics
Training Dynamics of ReLIFT. Figure 3 presents the training dynamics of rewards, response lengths, and the number of hardest questions during RL and ReLIFT training. As shown in Figure 3(a), both the RL and ReLIFT models exhibit an increasing trend in rewards, indicating progressive improvement in performance. Notably, the reward achieved by ReLIFT consistently surpasses that of standard RL, suggesting that ReLIFT is more effective in optimizing the target objective. Regarding response lengths, as depicted in Figure 3(b), RL tends to generate shorter responses as training progresses, which may reflect a diminishing capacity to address more challenging questions. In contrast, ReLIFT demonstrates a gradual increase in response length, implying greater computational effort and a higher potential to solve difficult problems. Furthermore, when examining the number of hardest questions encountered during training (Figure 3(c)), ReLIFT consistently faces fewer such instances compared to RL, and this number decreases over time.
Maintaining Exploration. As shown in Figure 3(d), the entropy of the RL model steadily declines throughout training, indicating reduced exploration. In contrast, the entropy of ReLIFT remains relatively high and exhibits persistent fluctuations, consistently surpassing that of RL. This sustained exploration allows ReLIFT to continue discovering novel solutions, thereby facilitating ongoing performance improvements even in the later stages of training. Actually, our method continues to exhibit a upward trend in reward over the final 500 training steps, underscoring the benefits of sustained exploration for long-term optimization.
3.4 Ablation Study and Analysis
@@
ReLIFT | 28.2 | 1666 | 20.1 | 1093 | 64.9 | 943 | 85.2 | 405 | 55.9 | 796 | 52.5 | 1463 | 51.1 | 1061 |
ReLIFT(all) | 7.8 | 26583 | 9.0 | 26704 | 31.6 | 22025 | 54.2 | 16984 | 26.4 | 21670 | 13.7 | 27938 | 23.8 | 23650 |
ReLIFT(uniform) | 23.2 | 3103 | 19.3 | 2762 | 62.3 | 1389 | 86.0 | 962 | 52.4 | 1426 | 51.4 | 655 | 49.1 | 1716 |
ReLIFT(random) | 22.1 | 1131 | 13.9 | 1527 | 59.5 | 449 | 84.2 | 120 | 44.7 | 433 | 48.4 | 22644 | 45.5 | 4367 |
We conduct an ablation study to assess the effectiveness of our ReLIFT approach. Specifically, we compare the following three ablation settings:
ReLIFT(all): Alternates one fine-tuning step after each RL step on the same batch.
ReLIFT(uniform): Applies fine-tuning at fixed, uniform intervals (every 8 RL steps), matching the number of demonstrations with ReLIFT.
ReLIFT(random): Performs interleaved FT. The buffer is populated based on the number of Hardest questions, but instead of using the hardest questions, it is randomly filled with other, non-hardest questions.
As summarized in Table 2, the results show that ReLIFT (all) quickly collapses, probably due to conflicting optimization objectives between RL and FT. ReLIFT (uniform) not only reduces accuracy but also leads to longer responses compared to ReLIFT. ReLIFT (random) causes a dramatic increase in average response length during training, while still underperforming ReLIFT. We also analyze the number of demonstrations of different methods and SFT, as depicted in Figure 4. ReLIFT requires only 13% of the demonstrations used in SFT.
Overall, these findings demonstrate that simply alternating RL and fine-tuning is ineffective; both the scheduling and the selection of fine-tuning training data are critical. ReLIFT applies fine-tuning more frequently in the early stages, when the model’s reasoning ability is relatively weak, and focuses on the hardest questions. This enables the model to efficiently acquire new knowledge without sacrificing performance or conciseness. In contrast, all, uniform, or random fine-tuning scheduling leads to suboptimal results and unnecessary inflation of response length.

### v2

@@ 3 Experimental Setup
Evaluation. For our evaluation, we primarily consider five widely adopted mathematical reasoning benchmarks: AIME 2024, AIME 2025, AMC Li et al. (2024), OlympiadBench He et al. (2024), and MATH500 Hendrycks et al. (2021). Due to the relatively small test sets for AIME 2024, AIME 2025, and AMC, we report avg@32 for these datasets. For OlympiadBench and MATH500, we use pass@1 as the evaluation metric. Since our reinforcement learning training is mainly centered on mathematical reasoning, we also assess generalization performance on MMLU-Pro Wang et al. (2024). To prevent data contamination, we shuffle the multiple-choice options. A testing temperature of 0.6 was used for all evaluations.
Baselines. We benchmark ReLIFT against the following baselines using Qwen2.5-Math-7B Yang et al. (2024) as the base model. For SFT methods, we fine-tuning the base model on the OpenR1-Math-46k-8192 dataset. For RL methods, we include GRPO Shao et al. (2024) trained on the same 46k dataset; SimpleRL-Zero Zeng et al. (2025), which applies GRPO to approximately 24k mathematical samples from GSM8K Cobbe et al. (2021) and MATH dataset Hendrycks et al. (2021); OpenReasoner-Zero Hu et al. (2025), a PPO-based approach trained on 129k multi-source samples; PRIME-Zero Cui et al. (2025), which conducts policy rollouts on 150k NuminaMath queries with implicit process rewards; and Oat-Zero Liu et al. (2025), which removes the standard deviation in GRPO advantage calculation, and is trained on the MATH dataset. For methods combining SFT and RL, we evaluate three approaches: RL w/ SFT loss, which incorporates a SFT loss term into the GRPO objective trained on the 46k dataset; LUFFY Yan et al. (2025), a mixed-policy GRPO approach that also utilizes the 46k dataset; and SFT then RL, a two-stage method that applies RL training after an initial SFT phase. Specifically, we evaluate two versions of SFT then RL: SFT then RL (v1) involves taking our trained SFT model and training it with RL for 300 steps on a held-out split of the OpenR1-Math-220k dataset, which contains about 49,000 prompts, similar to previous work Cui et al. (2025); Yan et al. (2025). For SFT then RL (v2), to fairly compare with ReLIFT, which uses 8,640 demonstration samples, we first perform SFT on a subset of 8,640 samples from OpenR1-Math-46k-8192. Then we conduct GRPO training on the full OpenR1-Math-46k-8192 dataset. For a more detailed setup for training these baselines, please refer to Appendix B.
4 Experimental Results
@@
| Qwen2.5-Math-7B | ||||||||||||||
| Qwen-Math | 16.4 | 2125 | 6.9 | 2039 | 45.5 | 1493 | 65.2 | 1171 | 30.1 | 1594 | 24.9 | 830 | 31.5 | 1542 |
| Qwen-Math-Instruct | 9.3 | 4065 | 8.2 | 3590 | 40.5 | 2905 | 78.0 | 1774 | 36.6 | 2888 | 34.0 | 4274 | 34.0 | 3249 |
| Previous RLVR methods | ||||||||||||||
| SimpleRL-Zero | 27.0 | - | 6.8 | - | 54.9 | - | 76.0 | - | 34.7 | - | 34.5 | - | 39.0 | - |
@@
| Replication of SFT and RL | ||||||||||||||
| SFT | 26.9 | 7344 | 25.5 | 7059 | 59.8 | 5675 | 85.6 | 3503 | 53.6 | 5662 | 44.4 | 3956 | 49.3 | 5533 |
| RL | 21.1 | 3287 | 17.5 | 2904 | 62.1 | 2127 | 85.8 | 1277 | 48.7 | 2178 | 46.3 | 1276 | 46.9 | 2175 |
| Methods combining SFT and RL | ||||||||||||||
| RL w/ SFT loss | 26.9 | 7344 | 23.1 | 7059 | 59.8 | 5675 | 85.6 | 3503 | 53.6 | 5662 | 44.4 | 3956 | 48.9 | 5508 |
| SFT then RL (v1) | 29.4 | 5581 | 21.0 | 5328 | 63.1 | 3908 | 87.8 | 2224 | 55.3 | 3772 | 47.6 | 2258 | 50.7 | 3845 |
| SFT then RL (v2) | 26.8 | 6453 | 23.0 | 6231 | 63.8 | 4586 | 88.4 | 2682 | 54.5 | 4615 | 48.9 | 2638 | 50.9 | 4534 |
| LUFFY | 27.3 | 5741 | 23.0 | 5267 | 63.5 | 3477 | 85.4 | 2361 | 53.6 | 3758 | 52.6 | 2241 | 50.9 | 3808 |
| ReLIFT | 28.3 | 5062 | 22.9 | 4571 | 65.1 | 3540 | 87.4 | 2241 | 57.8 | 3352 | 53.9 | 2248 | 52.6 | 3502 |
@@ 4.1 Main Results
| Model | GPU Hours | Demonstrations |
| SFT | ||
| RL | ||
| RL w/ SFT loss | ||
| SFT then RL v1 | ||
| SFT then RL v2 | ||
| LUFFY | ||
| ReLIFT |
SOTA performance with Qwen2.5-Math-7B. Table 1 presents our main results, comparing ReLIFT against four previous state-of-the-art RLVR methods and replication of SFT and RL, all based on the Qwen2.5-Math-7B model. Across five challenging competition-level reasoning benchmarks and one out-of-distribution benchmark, ReLIFT establishes a new state-of-the-art with an overall accuracy of 52.6%, surpassing all baselines. Notably, ReLIFT consistently achieves either the best or second-best performance on every individual benchmark, indicating ReLIFT achieves both higher reasoning performance and better generalization.
Comparing against methods combining SFT and RL. When compared with SFT and RL methods trained on the same dataset, ReLIFT outperforms all methods that combine RL and SFT, including RL w/SFT loss, SFT then RL, and LUFFY, as shown in Table 1. ReLIFT also produces more concise responses. Additionally, an analysis of the required demonstration data and GPU hours in Table 2 reveals that ReLIFT requires less training time and demonstration samples while achieving the best performance. These results underscore the effectiveness and efficiency of ReLIFT.
@@ 4.2 Training Dynamics
Figure 3 presents the training dynamics of rewards, response lengths, and training entropy during RL and ReLIFT training. The reward achieved by ReLIFT consistently surpasses that of RL, suggesting that ReLIFT is more effective in optimizing the target objective. Regarding response lengths, RL tends to generate shorter responses as training progresses, which may reflect a diminishing capacity to address challenging questions. In contrast, ReLIFT demonstrates a gradual increase in response length, implying a higher potential to solve difficult problems. The entropy of the RL model steadily declines throughout training, indicating reduced exploration. In contrast, the entropy of ReLIFT remains relatively high and exhibits persistent fluctuations, consistently surpassing that of RL. This sustained exploration allows ReLIFT to continue discovering novel solutions, thereby facilitating ongoing performance improvements even in the later stages of training.
@@
| Qwen2.5-Math-1.5B | ||||||||||||||
| Qwen-Math-1.5B | 2.8 | 1971 | 1.3 | 2058 | 19.5 | 1533 | 45.4 | 1108 | 18.5 | 1619 | 5.2 | 1982 | 15.5 | 1712 |
| Qwen-Math-1.5B-Instruct | 10.3 | 4042 | 7.6 | 3764 | 40.0 | 2809 | 77.8 | 1628 | 35.7 | 2834 | 33.7 | 4300 | 34.2 | 3230 |
| SFT | 12.7 | 7925 | 13.0 | 7719 | 40.9 | 6953 | 71.8 | 5986 | 33.8 | 7231 | 23.5 | 5405 | 32.6 | 6870 |
| RL | 9.8 | 2425 | 8.5 | 2195 | 44.1 | 1569 | 74.2 | 958 | 37.2 | 1695 | 31.4 | 1075 | 34.2 | 1653 |
| ReLIFT | 14.3 | 3691 | 10.0 | 3416 | 47.2 | 2207 | 76.4 | 1308 | 39.6 | 2274 | 31.7 | 1724 | 36.5 | 2437 |
| Qwen2.5-7B | ||||||||||||||
| Qwen-7B | 6.2 | 1632 | 2.4 | 1370 | 31.1 | 1076 | 63.8 | 617 | 26.5 | 1068 | 32.8 | 945 | 27.1 | 1118 |
| Qwen-7B-Instruct | 11.5 | 1939 | 6.1 | 1500 | 41.0 | 1629 | 73.0 | 1753 | 37.3 | 1550 | 53.1 | 3325 | 37.0 | 1949 |
| SFT | 15.7 | 7786 | 18.6 | 7533 | 49.8 | 6424 | 80.8 | 4252 | 44.3 | 6234 | 54.6 | 4869 | 44.0 | 6183 |
| RL | 15.5 | 1784 | 13.0 | 1487 | 50.4 | 1422 | 80.0 | 917 | 42.2 | 1438 | 57.6 | 954 | 43.1 | 1334 |
| ReLIFT | 19.1 | 5522 | 14.7 | 4831 | 51.9 | 3792 | 81.6 | 2211 | 46.1 | 3649 | 56.4 | 2173 | 45.0 | 3696 |
| LLaMA-3.1-8B | ||||||||||||||
| LLaMa-8B-Instruct | 5.6 | 1396 | 0.7 | 1401 | 20.7 | 1012 | 44.0 | 622 | 13.8 | 1063 | 38.1 | 357 | 20.5 | 975 |
| SFT | 0.8 | 2033 | 1.5 | 2032 | 11.5 | 2007 | 28.6 | 1903 | 8.6 | 1997 | 28.2 | 1558 | 13.2 | 1922 |
| RL | 1.8 | 940 | 0.0 | 909 | 10.8 | 797 | 28.2 | 616 | 7.3 | 790 | 39.4 | 549 | 14.6 | 767 |
| ReLIFT | 1.3 | 1236 | 0.2 | 1272 | 11.9 | 1048 | 35.2 | 741 | 11.0 | 1050 | 44.2 | 650 | 17.3 | 1000 |
4.3 Ablation Study and Analysis
We conduct an ablation study to understand what makes ReLIFT effective. Our findings reveal that the scheduling of fine-tuning and the selection of training data are crucial. Specifically, we compare the following three ablation settings: ReLIFT(all), which alternates one fine-tuning step after each RL step on the same batch, with the number of demonstrations matching SFT; ReLIFT(uniform), which applies fine-tuning at fixed, uniform intervals (every 8 RL steps), with the number of demonstrations matching ReLIFT; ReLIFT(random) which performs interleaved fine-tuning. The buffer is populated based on the number of Hardest questions, but instead of using the hardest questions, it is randomly filled with non-hardest questions.
As shown in Figure 5, ReLIFT (all) quickly collapses, probably due to conflicting optimization objectives between RL and FT. Similarly, using a uniform or random schedule for FT, as in ReLIFT(uniform) and ReLIFT(random), leads to lower accuracy and longer responses. In contrast, ReLIFT strategically applies fine-tuning more frequently in early training stages and focuses on the most challenging questions. This approach allows the model to efficiently gain new knowledge without sacrificing performance or conciseness. More detailed results are in Appendix D.
We also study the effect of the entropy coefficient, , on the fine-tuning step, using Qwen2.5-Math-7B. As seen in Figure 5, an of 0.0001 prove optimal, achieving the highest accuracy of 52.6. Varying in either direction from this value significantly decreased performance. These results highlight the importance of the entropy loss term for effectively blending SFT with RL. Therefore, we use consistently in our experiments.
@@ 4.4 Extension to More Models
To evaluate the generalizability of our method, we extend ReLIFT to three additional base models: the smaller Qwen2.5-Math-1.5B Yang et al. (2024), the weaker Qwen2.5-7B Team (2024), and Llama3.1-8b Dubey et al. (2024), which has a different architecture. As shown in Table 3, ReLIFT consistently achieves substantial improvements across all base models and six benchmarks, surpassing both SFT and RL. Our method consistently achieves a top-two ranking while maintaining an acceptable response length. These results underscore the broad applicability of ReLIFT in achieving superior performance and generalization across diverse model scales and architectures.
5 Related Works
5.1 RL-Driven Emergence of Reasoning Abilities
RL has recently emerged as a promising approach for enhancing the reasoning capabilities of LLMs. Early breakthroughs, such as OpenAI o-series Jaech et al. (2024); OpenAI (2025), DeepSeek-R1 Guo et al. (2025), and Kimi k-series Team et al. (2025b; a), have demonstrated that applying RLVR can lead to significant improvements in complex reasoning tasks. Notably, DeepSeek-R1 further shows that applying RLVR directly to a base model, without the need for intermediate SFT. Building on these foundations, Light-r1 Xie et al. (2025) incorporates curriculum learning to refine LLM’s reasoning skills, while Logic-rl Xie et al. (2025) adopts logic-driven reward functions to improve general reasoning ability. Deepscaler Luo et al. (2025) trains LLMs with progressively longer contexts as their performance improves. Based on GRPO, DAPO Yu et al. (2025) uses clip-higher to prevent entropy collapse, dynamic sampling for improved efficiency, and token-level policy gradient loss with overlong reward shaping to stabilize training.
5.2 Training Paradigms
Despite these significant advances, Yue et al. (2025) argue that RLVR does not enable LLMs to acquire genuinely new reasoning abilities beyond those present in their base models, as RLVR-trained models perform worse on pass@k than the base models at higher k values, indicating limited reasoning capability expansion. Similarly, Zhao et al. (2025) demonstrate that RL creates an ”echo chamber” effect, amplifying specific behavioral patterns already present in the pre-training data while suppressing others, thereby reinforcing existing behaviors rather than creating new reasoning capabilities. Furthermore, Cheng et al. (2025) finds that RL stifles exploration, causing models to converge on narrow behaviors and leading to performance plateaus on complex reasoning tasks. Research suggests that current RLVR methods do not fully leverage the potential of RL to create genuinely novel reasoning abilities in LLMs. In contrast, directly fine-tuning on demonstration data introduces new knowledge to LLMs. However, SFT requires high-quality, detailed demonstration data and tends to encourage memorization of the training data, often resulting in poor generalization to out-of-distribution scenarios Chu et al. (2025); Chen et al. (2025); Team et al. (2025b).
Some approaches attempt to combine SFT and RL. Typically, these methods first use SFT to stabilize the LLM’s ability and then apply RL to unlock more advanced reasoning skills in LLMs Guo et al. (2025); Wen et al. (2025). However, such combined methods remain limited: SFT depends significantly on the quality of demonstration data, and the subsequent RL phase is still confined to the knowledge present in the SFT-trained model, making it difficult for the model to acquire genuinely new information. LUFFY Yan et al. (2025) seeks to address these limitations by augmenting RL with off-policy reasoning traces, dynamically balancing imitation and exploration through a combination of off-policy demonstrations and on-policy rollouts during training. Nevertheless, this approach necessitates the preparation of off-policy data in advance and relies heavily on probability-sharpening techniques, which may restrict its applicability. In contrast, ReLIFT proposes a simpler and more direct method that integrates RL and SFT to leverage the strengths of both paradigms. Our approach encourages LLMs to explore when possible and integrate new information as needed, thereby pushing RL-driven reasoning beyond the boundaries of existing knowledge.
6 Conclusion
In summary, we systematically analyze the respective strengths of RL and SFT for LLM reasoning, finding that RL is more effective for learning easier questions while SFT is crucial for solving challenging questions. Building on these insights, we introduce ReLIFT, an approach that interleaves RL with online fine-tuning on the hardest questions. ReLIFT achieves state-of-the-art results with significantly less demonstration data, reduced training time, and more concise solutions. Future work will focus on scaling ReLIFT to larger models and developing more effective strategies for coordinating SFT and RL to further advance LLM reasoning and generalization.
Appendix A The Use of Large Language Models
We use LLMs to polish our writing. They act as a sophisticated assistant, improving the clarity, grammar, and style of our text.
Appendix B Main Experiment Details
RL Implementation. For RL, we omit the KL divergence loss term following Luo et al. (2025); Yan et al. (2025). Consistent with Dr.GRPO Liu et al. (2025) and LUFFY Yan et al. (2025), we exclude both length normalization and standard error normalization from the GRPO loss 1 in all experiments. The rollout batch size is set to 128, the number of rollout is 8, the update batch size is 64, and the learning rate is . A temperature of 1.0 is used during rollout generation for exploration. For Qwen2.5-Math-7B, We train 600 steps for all methods except SFT then RL, to be near with the GPU hour of ReLIFT. For other base models, we train 500 steps. All experiments are conducted on 8 x A800 GPUs.
SFT Implementation. For all our SFT models, we utilize the OpenR1-Math-46k-8192 dataset. Following a hyperparameter search on a sampled small-scale dataset, we evaluate batch sizes of 64, 128, and 256, in conjunction with learning rates of , and . The optimal configuration identified includes a batch size of 64 and a learning rate of . Employing the LLaMA-Factory framework Zheng et al. (2024), we train our models for 3 epochs.
RL w/ SFT Loss Implementation. We compute the on-policy loss on 7 on-policy samples and SFT loss on 1 off-policy sample per prompt. The remaining setup is consistent with our RL experiments.
SFT then RL Implementation. We conduct SFT using the same settings outlined in the SFT Implementation and train for 3 epochs accordingly. Subsequently, the RL phase follows the same configuration as described in RL Implementation, and we train the RL phase for 300 steps.
ReLIFT Implementation. During interleaved fine-tuning, we make use of demonstrations collected beforehand. The RL and fine-tuning configurations for ReLIFT are consistent with the settings previously described except learning rate. For both the RL and interleaved SFT phases, the learning rates are set to .
Evaluation. All evaluations are conducted using vllm Kwon et al. (2023), with the temperature set to 0.6 and the maximum number of tokens set to 8192. Results are verified using math-verify Face (2024). For the baselines, we independently reproduce and verify the results of all methods combining SFT and RL, such as LUFFY Yan et al. (2025), whereas the results for previous RLVR methods are reported as in the LUFFY paper.
Qwen2.5-Math Models. All evaluations are conducted using vllm Kwon et al. (2023), with a temperature setting of 0.6 and a maximum token limit of 8192. Results are validated using math-verify Face (2024). For baseline methods, we independently reproduce and verify results for all approaches combining SFT and RL, including LUFFY Yan et al. (2025). The results for other RL models are reported following the findings presented in the LUFFY paper.
Llama-3.1-8B. Because the full OpenR1-Math-46k-8192 dataset proved too challenging for Llama-3.1-8B, we created a smaller, more manageable dataset. We selected 11,000 samples from OpenR1 where DeepSeek-R1 provided the correct solution and the solution length was under 2,048 tokens. The SFT was then performed using a learning rate of , as suggested by Liu et al. (2025). The maximum response length is set to 2048 tokens for RL training and evaluation.
Motivation Experiment Details. We randomly sample a subset of 8,000 examples from the training dataset described in Section 3 as the training set, and select 1,000 examples as the validation set. RL and SFT are both conducted on this subset dataset for 120 steps, using the same parameters as in our main experiments.
Appendix C Chat Template
To encourage the base model to reason more effectively, following Yan et al. (2025), we adopt a complex chat template as shown in Figure 6. All our trained models, except LLaMA-3.1-8B, share the same system prompt for training and inference.
For LLaMA-3.1-8B, we do not use the above system prompt as we find the model cannot follow such an instruction. Thus, we use a simplified version as shown in Figure 7.
Appendix D Detailed Experiment Results